1. The following table are a random sample of 8 homes, shows the relationship between kilowatt-hours (thousands) used and the number of rooms in a private single-family residences. Homes Number of KWh (thousands) a) Compute the correlation coefficient b) Can we conclude that there is a positive association betwen KWh and number of in all private single-family residences? Use th 0.05 significance level. rooms c Develop the regression equation for electricity used(Wh) based d) Determine the coefficient of determination. e) Determine the 95% prediction interval of number of KWh, in thousands, for a six- on number of rooms. room houses (Standard error of estimate -0 4873 and y--10875
Answer #1

We need to find the sums ∑x , ∑y, ∑x*y, ∑x2 and ∑y2

We can use excel to find these sums :

F ry 30 16 35 9 20 12 20 25 197 25 16 25 9 16 9 16 25 14. 10 49 9 5 16 25 25 201 3 13 3 2669 C-y54534345 33 BX64735455 39 123456789 10

∑x = 39 , ∑y = 33 , ∑x*y =167, ∑x2 = 201 and ∑y2 = 141

r = rac{nsum (x*y)-sum (x)*sum (y)}{sqrt{[nsum x^2-(sum x)^2 ]*[nsum y^2-(sum y)^2]}}

r = (8 167) (39 33) 6 r20)-69- 30)1-16-141)-653 83)

r = 0.841209

b)

To check the significance of positive correlation we have to perform correlation test .

H0 : The significant positive correlation exists between X and Y

Ha : The significant positive correlation does not exists between X and Y

Test statistic : r = 0.8412

Critical value : We are given  α = 0.05, and we have n = number of pairs of x and y = 8

Since we are testing for positive correlation , this is one tailed test.

2-tailed testing 1-tailed testing 5 0.805 0.878 0.959 0.687 0.805 0.934 0.6080.729 0.882 0.669 0.7540.875 0.55 0.669 0.833 0.621 0.7070.8340.507 0.621 0.789 9 0.5820.666 0.798 0.472 0.582 0.750 10 0.5490.632 0.765 0.4430.549 0.715 0.685 12 0.497 0.5760.708 0.3980.497 0.658 0.476 0.5530.6840.380 0.476 0.634 14 0.458 0.5320.66 0.365 0.4580.612 0.441 0.5140.640.350.441 0.592 0.729 0.811 0917 0 6 0.521 0.602 0.735 0.419 0.521 13 15

Critical value = 0.621

Decision rule :

If correlation coefficient r is greater than critical value , there is significant positive correlation exists.

If correlation coefficient r is less than critical value , there is no significant positive correlation exists.

Here r = 0.8412 and critical value = 0.621

As r is greater than critical value , there is significant positive correlation exists between the X and Y

C) Regression equation :

kWh = b0 + b1 *number of rooms ; b0 is Intercept and b1 is Slope

b1 = rac{nsum (x*y)-sum (x)*sum (y)}{{[nsum x^2-(sum x)^2 ]}}

b1 = (8 167) - (39 33) (8 201) (39 *39)

b1 = 0.5632

b0 = ar{y}-(b_1*ar{x})

b0 = 4.125 - (0.5632*4.875)

b0 = 1.3793

kWh = 1.3793 + 0.5632*number of rooms

D) Coefficient of determination (r2) = r*r = 0.8412*0.8412

r2 = 0.7076

E) Prediction interval :

Lower bound = y_SE

Upper bound = hat{y} + SE

SE = 1-3 SST

We are given s = 0.4873 and ( x - ar{x} )2 = 10.875 , SSxx = sum x^2-rac{(sum x)^2}{n} = 6.125

To find t, we can use excel function =TINV(α, d.f)  

We are given confidence level = 0.95 , therefore α = 1 - 0.95 = 0.05

d.f = n - 2 = 8 - 2 = 6

=TINV(0.05,6) = 2.4469

SE = 1-3 SST = / - (2.4469 * О. 4873) *V1+ 110.875 8 6.125 + 25